# Unexpected fault activation in underground gas storage. Part II: Definition of safe operational bandwidths

### v1

@@ Abstract
Underground gas storage is a versatile tool for managing energy resources and addressing pressing environmental concerns. While natural gas is stored in geological formations since the beginning of the 20th century, hydrogen has recently been considered as a potential candidate toward a more flexible and sustainable energy infrastructure. Furthermore, these formations can also be used to sequester environmentally harmful gases such as CO2 securely. When such operations are implemented in faulted basins, however, safety concerns may arise due to the possible reactivation of pre-existing faults, which could result in (micro)-seismicity events. In the Netherlands, it has been recently noted that fault reactivation can occur “unexpectedly” during the life of an underground gas storage (UGS) site, even when stress conditions are not expected to cause a failure. The present two-part work aims to develop a modeling framework to investigate the physical mechanisms causing such occurrences and define a safe operational bandwidth for pore pressure variation for UGS operations in the faulted reservoirs of the Rotliegend formation, the Netherlands. In this follow-up paper, we investigate in detail the mechanisms and crucial factors that result in fault reactivation at various stages of a UGS project. The mathematical and numerical model described in Part 1 is used, also accounting for the effect of geochemical dissolution on reservoir and caprock weakening. The study investigates the risks of fault activation caused by the storage of different fluids for various purposes, such as long-term CO2 sequestration, CH4 and N2 injection and extraction cycles, and N2 permanent storage. The results show how geological configuration, geomechanical properties, and reservoir operating conditions may increase the risk of fault reactivation at various UGS stages. Furthermore, the analysis indicates that reservoir pressure near critical levels and reactivation during primary production may significantly contribute to potential fault instabilities. In light of these findings, operational guidelines for improving secure and effective storage operations are thereby presented.
@@
Fault reactivation , Underground Gas Storage , Hydrogen Underground Storage , Carbon Capture and Sequestration , Nitrogen Storage , Safety guidelines
journal: Geomechanics for Energy and the Environment
@@
[POLITO]organization=Department of Environment, Land and Infrastructure Engineering, Polytechnic of Turin, city=Turin, postalcode=10129, country=Italy
@@ 1 Introduction
The development of underground gas storage sites has been crucial in the global energy infrastructure since the early 20th century. The first operational underground storage system (USS) site dates back to 1915 (Al-Shafi et al., 2023). Over time, this technology has expanded with more than 600 facilities worldwide (Foulger et al., 2018) to effectively manage gas supply and demand on both seasonal and daily bases. While USS has typically focused on storing natural gas (UGS), it possesses the potential to store a variety of gases for diverse purposes, thus becoming a flexible tool for managing energy resources and addressing environmental concerns. This is especially important as Europe moves toward a net-zero greenhouse gas emission energy system (European Commission and Directorate-General for Climate Action, 2019). The scientific community has shown a growing interest in underground hydrogen storage (UHS) as an alternative to natural gas. Having the potential to accommodate significant volumes, ranging from tens of millions of cubic meters in caverns to potentially billions of cubic meters in depleted gas fields (Groenenberg et al., 2020), hydrogen versatility and absence of carbon emissions during production and utilization make it a more sustainable option in comparison to CH4. Additionally, carbon capture and storage (CCS) initiatives have identified underground storage of CO2 as a crucial component in mitigating greenhouse gas emissions and contrasting climate change.
@@
However, despite the widespread use of USS, inherent risks are associated with the potential reactivation of existing geological fractures (Ellsworth, 2013; Foulger et al., 2018; Keranen and Weingarten, 2018) and consequent induced seismicity. Although such events are statistically rare (Foulger et al., 2018), they demand significant attention due to their social and economic implications. In the Netherlands, natural gas (i.e., CH4) is currently stored in four UGS facilities: Bergermeer, Alkmaar, Norg, and Grijpskerk. They are located in the Carboniferous-Rotliegend formation in northern Europe, which is considered one of the most extensively explored petroleum systems worldwide (Gautier, 2003). Over the past few decades, significant seismic activity related to induced seismicity has been observed in this region (van Wees et al., 2014; Uta, 2017), particularly in three of the four UGS facilities (i.e., Bergermeer, Norg, and Grijpskerk). Recently, the Netherlands have also focused the attention on the feasibility of storing different gases in depleted gas fields and salt caverns as a crucial measure in the ongoing energy transition, by applying the same technology used in UGS to offshore CO2 storage through the Porthos project (https://www.porthosco2.nl), while research on hydrogen storage is in progress and primarily focused on salt caverns (UHS), although the potential for storage in abandoned gas fields is also being considered for the future. Additionally, nitrogen, which is utilized to convert high-calorific gas to low-calorific gas for residential heating and cooking purposes in Dutch households, is presently stored in a salt cavern near Heiligerlee (Ministry of Economic Affairs and Climate Policy, 2021; Muntendam-Bos et al., 2022).
@@
Given the future role that USS will play, it is crucial to understand why faults can become active and which mechanisms are potentially prone to lead to greater instability. During the cyclic or more permanent storage of these fluids, fault reactivation can occur when the natural stress regime on a fault surface is altered due to changes in pore pressure $P$ within a reservoir. The interaction between human activities at depth and the existing stress conditions determines the initiation, amount of slip, and extent of fault reactivation (Segall et al., 1994; Hettema et al., 2000; Candela et al., 2019). From a physical point of view, a fault remains locked if the applied shear stress does not exceed the frictional resistance to sliding, and this concept is well understood in terms of the effective stress principle, expressed by the equation $\tau<\tau_{L}=c+\mu(\sigma_{n}-P_{f})$ (Jaeger et al., 2007). Indeed, various factors, such as shear stress ($\tau$, $\tau_{L}$), cohesion ($c$), friction coefficient ($\mu$), normal compressive stress ($\sigma_{n}$), and fluid pressure ($P_{f}$), influence the fault behavior and susceptibility to seismic activity. Changes in these factors can occur over different timescales.
@@
On one hand, short-term processes can be interpreted through a geomechanical perspective involving injection/extraction activities lasting decades. In this context, most human-induced seismic events can be easily explained, as they occur due to fluid injection causing pressure to rise beyond the initial level, ultimately leading to the shear stress on the fracture surface to achieve its limit strength (Walsh and Zoback, 2016; Franceschini et al., 2024). However, a subset of recorded events cannot be accounted for by this mechanism. These events, identified as “unexpected” seismic events, occur when the pressure falls within the range typically experienced during primary production. In this scenario, fault reactivation occurs not only during primary production or gas storage at pressures exceeding the natural value $p_{i}$ (Deflandre et al., 2018; Cesca et al., 2014; Zhou et al., 2017), but also during cushion gas injection or producing and storing phases with pore pressures below the natural value but above a minimum pressure (Hager and Toksoz, 2009; Kraaijpoel et al., 2013; TNO, 2015; Nederlandse Aardolie Maatschappij BV, 2016). A typical time-behavior of the fluid pressure variations expected in UGS and UHS applications is sketched in Figure 1. The recorded seismic events are unexpected because the stress state should have already been experienced by the reservoir and surrounding faults during previous production phases. Unraveling the underlying mechanisms behind these “unexpected” seismic events is crucial for ensuring the safe and efficient operation of underground gas storage facilities.
@@
On the other side, long-term processes, spanning hundreds or even thousands of years, primarily involve fluid-rock interactions that can alter fault rock fabric, mineralogy, frictional properties, and cohesion. Presently, the understanding of the specific impacts of different gases on the mechanical properties of faults and rock is poorly understood, but the potential risk associated with their storage has to be considered. Indeed, the feasibility of the new technologies, in terms of performance, efficiency, and safety of the storage, greatly depends on fluid-rock interaction.
In this follow-up paper of Franceschini et al. (2024), we conduct an investigation into the phenomenon of “unexpected” seismic events in underground gas storage facilities. By understanding the factors contributing to such events, we can establish enhanced guidelines and best practices for operating underground gas facilities during different storage phases. We present a comprehensive study to re-assess the fundamental geomechanical causes of seismicity and specific risk factors, including the influence of geochemical dissolution effects on reservoir and caprock weakening. This work has three specific aims. The first one is to improve the understanding of the physics-based processes behind the induced seismicity during both cycling and permanent gas storage. The second aim involves investigating which are the roles of the Rotliegend reservoirs peculiarities, i.e., the graben-fault bounds, and the compartmentalization into offset blocks by intra-field faults on favoring fault reactivation. Lastly, the third aim is to propose, based on the modeling outcomes, valuable insights into the development of general guidelines for safe operational bandwidths for reservoir storage conditions during pressure cycling (and noncycling) of underground storage of CH4, CO2, H2, and N2 in depleted gas fields. These guidelines are intended to serve as a valuable tool, contributing to the effectiveness of risk assessment and mitigation strategies in underground gas storage operations. A few preliminary outcomes regarding UGS are already reported in Teatini et al. (2019, 2020) and Franceschini et al. (2024).
@@
The analysis that follows is accomplished by using the mathematical and numerical framework developed in Part I of this work (Franceschini et al., 2024), where a Finite Element (FE) geomechanical simulator is implemented with a Lagrange multiplier-based treatment of the frictional contact conditions for reproducing the fault behavior. The simulator is supplemented also with visco-elasto-plastic constitutive laws, e.g. Isotton et al. (2019), and fault activation is ruled by the Coulomb frictional criterion. Pressure change within the faults, variation of Coulomb’s parameters due to slip-weakening, and the rheology of the caprock are properly accounted for. Because of the main aims listed above, the modeling study is not focused on a specific UGS field, but on a conceptual reservoir and fault system that realistically represents the main geologic features of the Rotliegend formation system.
@@
The paper is organized as follows. The typical setting of the Rotliegend reservoirs and a few examples of seismicity recorded during UGS are initially presented. Then, fluids under investigation are described, and the modeling set-up is outlined. The various scenarios are then developed to understand the geomechanical behavior of the faulted system, first analyzing the parameters involved in UGS and then comparing the scenarios using different fluids. Modeling results are presented, and the mechanisms responsible for fault reactivation during the different phases are identified. The modeling outcomes are discussed, highlighting the peculiarities of the storage scenarios with respect to induced seismicity during primary production as addressed in previous studies, ranking the natural features and anthropogenic factors more prone to cause fault reactivation, and illustrating general operational guidelines to reduce the probability of reactivation occurrences.
@@
Figure 1: Sketch of the main stages during the reservoir lifespan in terms of pressure change over time for CH4, CO2, H2 and N2: Primary Production (PP) for CH4 production, Storage (ST) for CO2 and N2 storage, Cushion Gas Injection (CGI) for H2 and CH4 injection, and Underground H2 Storage (UHS) and Underground Gas Storage (UGS) cycles for seasonal pressure fluctuations during CH4/H2 storage (UGS-S/UHS-S) and production (UGS-P/UHS-P).
2 Induced seismicity in Rotliegend UGS reservoirs
@@
The Rotliegend formation stretches over a large region covering north-western Germany, the Netherlands, with an off-shore portion that extends and includes part of the UK. The Rotliegend sandstones have been attributed to deposition in a variety of arid, terrestrial environments, among which ephemeral fluvial (wadi) systems, various types of eolian deposits, desert-lake environments, and adjacent sabkhas are dominant. The sealing nature of Zechstein salt is the main reason for the existence of the large amount of traps that sealed gas migrated from the underlying coals and carbonaceous shales of the Carboniferous Coal Measures for approximately 150 million years (Gautier, 2003). The Rotliegend unit spans a depth range between 2000 and 4700 m and it is generally made of permeable sandstone (average 100-200 mD), with high net-to-gross and porosity (average 15-20%). The thickness of the reservoir sands is generally over 50 m and rarely exceeds 300 m. Faulting is the peculiarity of the geologic unit, with the majority of the faults that are Triassic-Tertiary features nucleated on Mid Palaeozoic fault patterns and reworked during the Variscan orogeny. The Mesozoic rifting phase established the general extensional normal fault pattern (De Jager, 2007).
As a consequence, many fields are formed by a number of rhomboid-shaped dipping fault blocks, positioned one next to each other. The trapping mechanism is structural, by the juxtaposition of the Rotliegend reservoir blocks against the thick evaoporitic Zechstein formation. Moreover, the hydraulic connectivity between the various blocks is extremely poor, causing a clear compartmentalization of the reservoirs. Fault seal and, secondarily, stratigraphic barriers and diagenesis, are the causes of this compartmentalization (van Hulten, 2010). The leftomost frame in Figure 2 sketches as a representative examples the structural map of the Norg UGS reservoir.
Refer to caption Refer to caption
Figure 2: Left: Map of the Norg UGS site (in blue) with traces of the bounding faults (black lines) and localization of the recorded seismic events. Right: pressure evolution in time at the Grijpskerk UGS field. The location of the two seismic events in 1997 and 2015 are reported (modified from TNO report (2015)).
The reservoir is bounded by normal faults with a significant (up to a 250 m) throw and consist of a few compartments also separated by internal faults. The gas fields are 2500-3500 m deep, with the Rotliegend reservoir rock characterized by an average net thickness of 150-200 m. As usual for UGS plants, the reservoir is a former partially-depleted gas field that was converted to underground gas storage after a primary production (PP) period. PP is generally followed by a cushion gas injection (CGI) phase when fluid pressure in the reservoir is almost risen to the original value, and then by the UGS injection and production cycles. The rightmost frame of Figure  2 summarizes the evolution of the average pressure in Grijpskerk, another representative UGS field. It must be highlighted that pressure differences over reservoir blocks can be as high as about 7 MPa. Gas is injected at temperatures similar or slightly below those at the reservoir level, thus minimizing thermal stresses.
The timing and location of the seismic events recorded in the vicinity of the UGS reservoirs is also shown in Figure 2. In Norg, the time of the first earthquake coincided with low pressure level during PP. The second seismic event occurred at the end of the CGI phase, with the pressure in the reservoir close to the original pressure level. In Grijpskerk, a first seismic event took place at the end of PP in 1997, and a second one much later, in 2015, almost at the end of a UGS production phase (TNO, 2015). Four seismic events with a magnitude $M\in[3.0,3.5]$ were detected in 1994 and 2001 during PP in the Bergermeer reservoir. The events were located at the tip of the central fault separating the two main blocks (TNO, 2015). A down-hole micro-seismic array characterized by a magnitude of completeness below $M=0$ was established during the CGI phase. The array recorded a large number of micro-seismic events ($M<0$) on the faults at the reservoir depth and a main event with $M=0.9$ along the central fault when the pressure difference between the two blocks peaked about 4 MPa at the beginning of 2013 (Baisch et al., 2016).
3 Chemo-mechanical effects by CO2, H2 and N2 on rocks and faults
3.1 Effect on rock and caprock

### v4

# Unexpected fault activation due to underground gas storage in produced reservoirs. Part II: Definition of safe operational bandwidths

@@ Abstract
Underground gas storage is a versatile tool for managing energy resources and addressing pressing environmental concerns. While natural gas is stored in geological formations since the early 20th century, hydrogen has recently been considered as a potential candidate toward a more flexible and sustainable energy infrastructure. Furthermore, these formations can additionally capture gases that contribute to climate change, such as CO2. When such operations are implemented in faulted basins, however, safety concerns may arise due to the potential reactivation of pre-existing faults, which could trigger (micro)-seismicity events. In the Netherlands, it has been recently noted that fault reactivation can occur “unexpectedly” during the life of an underground gas storage (UGS) site, even when stress conditions are not expected to cause a failure. The present two-part work aims to develop a modeling framework to investigate the physical mechanisms causing such occurrences in previously produced gas reservoirs and define a safe operational bandwidth for pore pressure variation for UGS operations in the faulted reservoirs of the Upper Rotliegend Group, the Netherlands. This follow-up paper investigates in detail the mechanisms and crucial factors that result in fault reactivation at various stages of a UGS. The mathematical and numerical model described in Part I is used, also considering how the presence of stored gases may influence the mechanical properties of the reservoir and caprock, in particular the Young modulus. The study investigates the hazard of fault activation caused by the storage of different fluids for various purposes, such as long-term CO2 sequestration, CH4 and H2 injection and extraction cycles, and N2 injection as cushion gas. The results show how geological configuration, geomechanical properties, and reservoir operating conditions may increase the hazard of fault reactivation at various UGS stages. Furthermore, the analysis indicates that reservoir pressure near critical levels and reactivation during primary production may significantly contribute to potential fault instabilities. Operational guidelines for improving secure and effective storage operations are thereby presented.
@@
Carbon Capture and Sequestration , Critical pressure , Fault reactivation , Safety guidelines
journal: Netherlands Journal of Geosciences
@@
[POLITO]organization=Department of Environment, Land and Infrastructure Engineering, Politecnico di Torino, city=Turin, postalcode=10129, country=Italy
@@ 1  Introduction
The development of underground gas storage sites (UGS) has been crucial in the global energy infrastructure since the early 20th century. The first operational underground storage system (USS) site dates back to 1915 in the Welland gas field (Canada) (Al-Shafi et al., 2023) . Over time, this technology has expanded with more than 600 facilities worldwide (Foulger et al., 2018) to effectively manage gas supply and demand on both seasonal and daily bases (Verga, 2018). While USS has typically focused on storing natural gas (i.e., mainly CH4) through UGS, it possesses the potential to store a variety of gases for diverse purposes, thus becoming a flexible tool for managing energy resources and addressing environmental concerns. This is especially important as Europe moves toward a net-zero greenhouse gas emission energy system (European Commission and Directorate-General for Climate Action, 2019). The scientific community has shown a growing interest in underground hydrogen storage (UHS) as an alternative to natural gas. Having the potential to accommodate significant volumes, ranging from tens of millions of cubic meters in caverns to potentially billions of cubic meters in depleted gas fields (Groenenberg et al., 2020), hydrogen versatility and absence of carbon emissions during production and utilization make it a sustainable option. Additionally, carbon capture and storage (CCS) initiatives have identified underground storage of CO2 as a crucial component in mitigating greenhouse gas emissions and contrasting climate change.
However, despite the widespread use of USS, one of the important aspects to be considered is the potential reactivation of existing faults and consequent induced seismicity (Ellsworth, 2013; Foulger et al., 2018; Keranen and Weingarten, 2018). Although such events are statistically rare (Foulger et al., 2018), in terms of the number of affected sites rather than the frequency of events within a single reservoir, they demand significant attention due to their social and economic implications (Van der Voort and Vanclay, 2015). In the Netherlands, natural gas is currently stored in four UGS facilities: Bergermeer, Alkmaar, Norg, and Grijpskerk. They are located in the Upper Rotliegend group that widely extends in Central Europe and is considered one of the most extensively explored petroleum systems worldwide (Gautier, 2003). Over the past few decades, seismic activity associated to USS has been observed in this region (Van Wees et al., 2014; Uta, 2017), particularly in three of the four UGS facilities (i.e., Bergermeer, Norg, and Grijpskerk). Aside from the Castor in Spain (Cesca et al., 2014; Vilarrasa et al., 2021) and Hutubi in China (Jiang et al., 2021; Liu et al., 2023), these Dutch reservoir represent the only cases where UGS activities have been definitively associated to induced seismicity events (Vilarrasa et al., 2021).
@@
Recently, the Netherlands have also focused attention on the feasibility of storing different gases in depleted gas fields and salt caverns as a crucial measure in the ongoing energy transition. The same technology used in UGS is being applied to offshore CO2 storage through the Porthos project (Porthos CO2 Transport and Storage C.V., 2024), while research on hydrogen storage is in progress and primarily focused on salt caverns, although the potential for storage in abandoned gas fields is also being considered for the future. Additionally, besides being used as cushion gas (Shoushtari et al., 2023), nitrogen is utilized to convert high-calorific gas to low-calorific gas for residential heating and cooking purposes in Dutch households. Currently, nitrogen is stored in a salt cavern near Heiligerlee for this purpose (Ministry of Economic Affairs and Climate Policy, 2021; Muntendam-Bos et al., 2022).
@@
Given the future role that USS will play, it is crucial to understand why faults can become active and which mechanisms are potentially prone to cause instability. During the cyclic or permanent storage of these fluids, activation can occur when the natural stress regime on a fault surface is altered due to changes in pore pressure within a reservoir, although additional mechanisms may also contribute. The combined effects of human-induced pressure and stress changes at depth, the pre-existing stress field, and the frictional and mechanical properties of the rocks and faults determine the initiation, amount of slip, and extent of reactivation (Segall et al., 1994; Hettema et al., 2000; Candela et al., 2019).
@@
Over a short- to mid-term timeframe (e.g., from days to years), fault reactivation can be interpreted through a geomechanical approach involving pressure variations following injection/extraction activities that can last for decades. In this context, most human-induced seismic events can be easily explained, as they are triggered by fluid injection that increases pore pressure beyond its initial reservoir pressure, thereby driving shear stress on the fault surface to its failure limit (Walsh and Zoback, 2016; Franceschini et al., 2024). However, a subset of recorded events cannot be accounted for by this mechanism. These events, referred to as “unexpected” seismic events, typically exhibit small magnitudes and occur when the pressure falls within the pressure range already experienced during primary production, that is, pressures lower than the initial value $P_{i}$ but still above the minimum pressure generally experienced by the reservoir at the end of depletion $P_{\min}$. In this scenario, fault reactivation occurs not only during primary production or gas storage at pressures exceeding the initial value $P_{i}$ (Deflandre et al., 2018; Cesca et al., 2014; Zhou et al., 2017), but also during injection or producing and storing phases in which pore pressures remain within the previously experienced pressure interval (Hager and Toksoz, 2009; Kraaijpoel et al., 2013; TNO, 2015; Nederlandse Aardolie Maatschappij BV, 2016). A typical time-behavior of the fluid pressure variations expected in USS applications is sketched in Figure 1. The recorded seismic events in UGS are unexpected because they occur at a stress state already experienced by the reservoir and surrounding faults during previous production phases. However, the mechanical response of the reservoir–fault system may evolve over time, meaning that previously experienced stress conditions do not necessarily lead to identical system behavior (e.g. Pijnenburg et al., 2019b). Unraveling the underlying mechanisms behind these “unexpected” seismic events is crucial for ensuring the safe and efficient operation of underground gas storage facilities.
@@
On the other hand, long-term processes, spanning hundreds or even thousands of years, primarily involve fluid-rock interactions that can alter fault rock fabric, mineralogy, frictional properties, and cohesion. Currently, the understanding of the specific impacts of different gases on the mechanical properties of faults and rock is still poorly understood, but the potential risk associated with their storage has to be considered. Indeed, the performance, efficiency, and safety of any storage greatly depends on fluid-rock interaction.
In this follow-up paper of Franceschini et al. (2024), we conduct an investigation of the phenomenon of “unexpected” seismic events in underground gas storage facilities. By understanding the factors contributing to such events, we can establish enhanced guidelines and best practices for operating underground gas facilities during different storage phases. We present a comprehensive study to reassess the fundamental geomechanical causes of induced seismicity and specific risk factors, including the potential impact of gas type on reservoir and caprock mechanical weakening, particularly with regard to changes in elastic properties.
This work has four specific aims. The first is to improve the understanding of the physics-based processes responsible for induced seismicity during both cyclic and permanent gas storage. To this end, the analysis does not attempt to reproduce individual field events, but rather to identify the mechanical conditions and stress-path evolutions under which fault reactivation may occur during storage operations. The second aim is to investigate how the peculiarities of the Rotliegend reservoirs, such as graben-bounding faults and intra-field faults responsible for reservoir compartmentalization, affect fault reactivation. The third aim is to assess how different injected or stored fluids (CH4, CO2, H2, and N2), through their associated effects on rock elastic properties such as the Young’s modulus, impact the physical mechanisms of fault reactivation. The final aim is to provide, based on the modeling outcomes, recommendations for operational guidelines defining safe pressure bandwidths for reservoir storage conditions. These guidelines are relevant to pressure cycling (and noncycling) during the underground storage of CH4, CO2, H2, and N2 in depleted gas fields with configurations similar to those of the reservoirs located in the Rotliegend Group, and are intended to assist risk assessment and mitigation strategies in underground gas storage operations. A few preliminary outcomes regarding UGS are already reported in Teatini et al. (2019, 2020) and Franceschini et al. (2024).
It is important to emphasize that this study specifically addresses the hazard of fault reactivation in UGS sites that are hosted by depleted hydrocarbon reservoirs, where the mechanical behavior of the formation is already well constrained. Gas storage in other subsurface systems, such as saline aquifers, salt caverns and unexploited reservoirs, may differ substantially in terms of knowledge of geomechanical setting, and are not addressed by this work. As such, the findings and modeling framework presented here are not intended to be generalized to those settings.
@@
The analysis that follows is accomplished by using the mathematical and numerical framework developed in Part I of this work (Franceschini et al., 2024), where a Finite Element (FE) geomechanical simulator is implemented with a Lagrange multiplier-based treatment of the frictional contact conditions for reproducing the fault behavior. The simulator is supplemented with visco-elasto-plastic constitutive laws Isotton et al. (2019), and fault activation is ruled by the Coulomb frictional criterion. Pressure change within the faults, variation of Coulomb’s parameters due to slip-weakening, and the rheology of the caprock are properly accounted for. Because of the main aims listed above, the modeling study is not focused on a specific UGS field, but on a conceptual reservoir and on a fault system that realistically represents the main geologic features of the Upper Rotliegend group.
@@
The paper is organized as follows. The typical setting of the Rotliegend reservoirs and a few examples of seismicity recorded during UGS are initially presented. Then, fluids under investigation are described, and the modeling set-up is outlined. The various scenarios are then developed to understand the geomechanical behavior of the faulted system, first analyzing the parameters involved in UGS and then comparing the scenarios using different fluids. Modeling results are presented, and the mechanisms responsible for fault reactivation during the different phases are identified. The modeling outcomes are discussed, highlighting the peculiarities of the storage scenarios with respect to induced seismicity during primary production as addressed in previous studies, ranking the natural features and anthropogenic factors prone to cause fault reactivation, and illustrating general operational guidelines to reduce the probability of reactivation occurrences.
@@
Figure 1: Sketch of the main stages during the reservoir lifespan in terms of pressure change over time for CH4, CO2, H2 and N2: Primary Production ($P$) for CH4 production, Storage (ST) for CO2 and N2 storage, Cushion Gas Injection (CGI) for H2 and CH4 injection, and Underground H2 Storage (UHS) and Underground Gas Storage (UGS) cycles for seasonal pressure fluctuations during CH4/H2 injection and withdrawal. Notice that, according to the Netherlands regulation, $P$ must remain below $P_{i}$ regardless of the stage under consideration (Ministerie van Economische Zaken en Klimaat, 2022; TNO, 2018)
1.1 Induced seismicity in Rotliegend UGS reservoirs
@@
The Rotliegend Group extends over a wide region that includes North-Western Germany and the Netherlands, with an offshore portion into the southern North Sea and parts of the UK (Figure 2 a). The Rotliegend Group sandstones were deposited in a variety of arid, terrestrial environments, among which ephemeral fluvial (wadi) systems, various types of eolian deposits, desert-lake environments, and adjacent sabkhas were dominant. The sealing nature of the Zechstein Group salt is the main reason for the existence of the large number of traps that retained gas migrated from the underlying coals and carbonaceous shales of the Carboniferous Limburg Group for approximately 150 million years (Gautier, 2003). In the Dutch sector, the reservoirs of the Upper Rotliegend Group spans a depth range between 2,000 and 4,700 m, and is generally composed of permeable sandstones (average 100-200 mD), with high net-to-gross and porosity (average 15-20%). The thickness of the reservoir sands is generally over 50 m and rarely exceeds 300 m. Faulting is a characteristic feature of this geological unit. Most present-day faults represent reactivations of inherited Mid-Palaeozoic basement structures that were deformed during the Variscan orogeny and subsequently reactivated during Permian, Mesozoic, and Cenozoic tectonic phases (Geluk, 2005). During the Mesozoic, successive rifting episodes further reactivated and enhanced the extensional normal-fault pattern that characterizes the region (Ziegler and others, 1990; De Jager, 2007). As a consequence, many fields are formed by a number of rhomboid-shaped dipping fault blocks, positioned one next to each other. The trapping mechanism of gas is structural, by the juxtaposition of the Rotliegend reservoir blocks against the thick Zechstein formation. Moreover, in several Rotliegend fields, the hydraulic connectivity between fault-bounded blocks is limited, causing a clear compartmentalization of the reservoirs. This behaviour results from a combination of fault sealing, fault offsets that produce structural barriers through lithological juxtaposition, and stratigraphic and diagenetic heterogeneities (Van Hulten, 2010). Nevertheless, some Dutch Rotliegend reservoirs (e.g., Groningen) do contain transmissive faults (De Jager and Visser, 2017). In such settings, compartmentalization is mainly structural rather than hydraulic. Figure 2 c sketches a representative example of the structural map of the Norg UGS reservoir.
Figure 2: Key geological and seismological features of the Rotliegend formation. a): Tectonic provinces of the Carboniferous–Rotliegend Total Petroleum System with the hydrocarbon field centerpoints distinguished between oil and gas fields (modified after Gautier (2003)). b): Top view of the Norg UGS site (in green), with bounding fault traces (red lines) and recorded seismic events (blue circles). The dashed purple line surrounding Blocks 2a, 2b, and 2c illustrates an example of reservoir compartmentalization (modified after TNO (2015)). c): Interpreted seismic profile across the northeast Netherlands, showing the Norg gas field located in fault-bounded blocks at Rotliegend level (modified after de Jager and Geluk (2007)). d)t: Average pressure evolution at the Norg UGS reservoir over time, with annotated seismic events from 1993 and 1999 (modified after Nederlandse Aardolie Maatschappij BV (2016)).
The reservoir is bounded by normal faults with a significant throw (up to a 250 m) and consist of a few compartments that are separated by internal faults. The gas fields are 2,500-3,500 m deep, with the Rotliegend reservoir rock characterized by an average net thickness of 150-200 m (Figure 2 c. As usual for UGS plants, the reservoir is a partially-depleted gas field that was converted into an UGS after a primary production (PP) period. PP is generally followed by a quick gas injection (CGI) phase when fluid pressure in the reservoir is risen to almost the original value, and then by the UGS injection and withdrawal cycles. Gas is injected at temperatures similar to, or slightly below, those at reservoir depth, thus minimizing thermal stresses. Figure 2 d summarizes the evolution of the average pressure in Norg. The timing and location of the seismic events recorded in the vicinity of the UGS reservoir is also shown in Figure 2 b,d. In Norg, the time of the first earthquake ($M=1.5$) coincided with low pressure level during PP. The second seismic event $M=1.1$ occurred at the end of the CGI phase, with the pressure in the reservoir close to the original pressure level. In Grijpskerk, where pressure differences between reservoir blocks can be as high as about 7 MPa, a first seismic event occurred at the end of PP in 1997 at a pressure about 11 MPa below the initial value, and a second one in 2015, near the end of a UGS production phase, at about 12 MPa below the initial conditions (TNO, 2015). Four seismic events with magnitudes $M\in[3.0,3.5]$ were detected in 1994 and 2001 during PP in the Bergermeer reservoir, when reservoir pressures had decreased by about 18–20 MPa from the initial pressure conditions. The events were located at the tip of the central fault separating the two main blocks (TNO, 2015). A down-hole microseismic array characterized by a magnitude of completeness below $M=0$ was established during the CGI phase. The array recorded a large number of microseismic events ($M<0$) on the faults at the reservoir depth and a main event with $M\approx 0.8$ along the central fault when the pressure difference between the two blocks peaked at about 4 MPa in early 2013 (Baisch et al., 2016). Across these Dutch Rotliegend UGS sites, injection-related events are of very small magnitude (typically $M<1.5$), and thus contribute only a negligible fraction of the cumulative seismic moment compared to the $M$ 33.53-3.5 events produced during primary depletion. Table 1 presents an overview of induced seismicity cases across the different reservoir sites, detailing the operational phase, activated fault geometry, and event magnitudes. Table 2 follows with a summary of the natural stress regimes, fault orientations, and key geological and operational characteristics of the reservoirs. Taken together, these cases show a common pattern: fault reactivation can occur not only at high pore pressures but also within pressure ranges that were previously experienced without seismicity. This indicates that the key control on stability is the evolving stress path of the reservoir–fault system, influenced by differential compaction, fault-block displacement, and changes in fault-zone hydraulic properties. These shared features form the basis of our geomechanical modelling, which aims to investigate the physical mechanisms capable of triggering such “unexpected” seismicity during UGS operations.
Reservoir Type / gas Activated fault geometry $\Delta P$ at event (MPa) Magnitude
$\leq$1 1–3 $\geq$3
Norg Nederlandse Aardolie Maatschappij BV (2016); TNO (2015)      The Netherlands UGS (CH4) Slip motion on a sub-vertical NW–SE fault (PP) and a sub-vertical compressive SW–NE fault (UGS) PP (1993) $\sim-9$ UGS (1999) $\sim-2$ x (2)
Grijpskerk TNO (2015)      The Netherlands UGS (CH4) Large uncertainty in locating the active fault(s) UGS (1997) $\sim-11$ UGS (2015) $\sim-12$ x (2)
Bergermeer TNO (2015); Orlic et al. (2013)      The Netherlands UGS (CH4) Central fault with strike SSW and dip 60–65 PP (1994): $\sim-18$ PP (2001): $\sim-20$ CG(2013): $\sim-15$ x (1–CG) x (1–PP)
Table 1: Inventory of induced seismicity cases in the Netherlands, categorized by activated fault geometry, pressure conditions, and event magnitude. $P$ = primary production, CG = cushion gas injection, UGS = underground gas storage. $\Delta P$ denotes the pressure difference at the time of the event with respect to the initial reservoir pressure $P_{i}$. The notation “x (M–Phase)” indicates a recorded seismic event of approximate magnitude $M$ during the specified operational phase.
Reservoir Stress regime Reservoir info Fault orientation
Norg Nederlandse Aardolie Maatschappij BV (2016); TNO (2015) Normal faulting; $\sigma_{v}>\sigma_{H}>\sigma_{h}$; SHmax $\approx$ 1.92 bar/10m (171° E of N), $\sigma_{h}$ $\approx$ 1.44 bar/10 m, and $\sigma_{v}$ $\approx$ 2.2 bar/10 m Depth: 2670 m; Thickness: 140 m. Primary production started in 1983 until 1995 with a max DP of about 117 bar; UGS started in 1995–1997 with a DP for cycle of about 57 bar Steeply dipping NW–SE and SW–NE; mostly non-sealing
Grijpskerk TNO (2015) Normal faulting; $\sigma_{v}>\sigma_{H}>\sigma_{h}$; SHmax oriented $\approx$ N170°E Depth: 3300 m; Thickness: 220 m. Primary production started in 1993 until 1994 with a max DP of about 53 bar; UGS started in 1997 with a DP for cycle of about 127 bar Steeply dipping NW–SE and NE–SW; mostly non-sealing
Bergermeer TNO (2015); Orlic et al. (2013) $\sigma_{v}$ = 25 MPa; $\sigma_{h\,min}$ = 9 MPa (effective stress); initial direction of the minimum horizontal stress is oriented NE–SW Depth: 2200 m; Thickness: 200 m. Primary production started in 1971 until 2005 with a max DP of about 213 bar; no activity between 2007 and 2010; cushion gas injection between 2010 and 2012; UGS started in 2013 Not specified here; typically NW–SE in region
Table 2: Overview of induced seismicity cases in relation to the stress regime, seismic monitoring, reservoir characteristics, and production history.
1.2 Chemo-mechanical effects by CO2, H2 and N2 on rocks and faults
1.2.1 Effect on CO2